A pipe has a cross sectional area of 50cmsq. Water is flowing through the pipe at a speed of 25 cm/sec. The pipe is used to fill up a cuboidal tank whose inner dimensions are 1m* 0.8m*0.2m. How many minutes will it take for the tank to be completely full?
Answer by rothauserc(4718)

The flow rate is equal to flow area times the flow velocity
:
we are given the flow area = 50cm^2 and the flow velocity = 25 cm/sec, therefore
the flow rate = 50cm^2 * 25 cm/sec = 1250cm^3/sec
:
Convert the dimensions of the cuboidal tank to cm
1m * 0.8m * 0.2m = 100cm * 80cm * 20cm
:
Calculate the volume of the cuboidal tank in cm
Vcm = 100 * 80 * 20 = 160000cm^3
:
now divide Vcm by the flow rate to determine how many seconds to fill the tank
:
(160000cm^3/sec) / 1250cm^3 = 128 seconds
:
now divide 128 seconds by 60 seconds in a minute to determine how many minutes to fill the tank
:
128 / 60 = 2.133333333 minutes approx 2.1 minutes